OpenStudy (alexh107):

How to do (1/4)^(-1/2)?

OpenStudy (mrs.ambrose614):

Hold on I will help

OpenStudy (alexh107):

Ok thanks

OpenStudy (mrs.ambrose614):

Okay first you have to simplify the expression

OpenStudy (alexh107):

How would I do that?

OpenStudy (mrs.ambrose614):

Your awnser would be -1/8

OpenStudy (alexh107):

The answer in my book says 2

OpenStudy (mrs.ambrose614):

Then thatwould be your awnser

OpenStudy (alexh107):

I have to show my work though and I'm not sure how to get it

OpenStudy (mrs.ambrose614):

For the work just put you have to simplify the expression by breaking it down

OpenStudy (s4sensitiveandshy):

\[x^{-a} = \frac{1}{x^a}\]

OpenStudy (alexh107):

I know how to do negative exponents it's just the fraction part is confusing me

OpenStudy (s4sensitiveandshy):

oh okay. you have to convert exponent to a radical

OpenStudy (s4sensitiveandshy):

\[x^{\frac{1}{2}} =\sqrt{x}\]

OpenStudy (alexh107):

Oh okay. I think I get it. Would you make it sq rt 4 and then get 2

OpenStudy (s4sensitiveandshy):

\[\large x^{\frac{m}{n}} = \sqrt[n]{x^m}\]

OpenStudy (s4sensitiveandshy):

well, first of all how would you rewrite the expression with positive exponent?

OpenStudy (alexh107):

4^1/2?

OpenStudy (s4sensitiveandshy):

remember: the entire fraction raised to a negative exponent

OpenStudy (s4sensitiveandshy):

\[\frac{ 1 }{ \sqrt{\frac{1}{4}} }\] \[\frac{ 1 }{(4)^ \frac{1}{2} } \rightarrow 1 \times \frac{ 4^\frac{ 1 }{ 2 } }{ 1 } \rightarrow \sqrt{4}=2\] your answer is correct: make sure 1/ (1/4)^(1/2)

OpenStudy (alexh107):

Oh okay. That makes sense. Thanks for the help.

OpenStudy (s4sensitiveandshy):

correction: \[\frac{ 1 }{ (\frac{ 1 }{ 4 } )^\frac{ 1 }{ 2 }}\]

OpenStudy (s4sensitiveandshy):

you're welcome. good job.